A compound is made up of silver and oxygen. The sample itself weighs 1.7734 g and you have isolated 1.6510 grams of silver. What is the molar mass, grams in the sample, moles in the sample, molar ratio of Ag: O, empirical formula of the compound, molar mass of the empirical formula ( for Ag and O)
1
Expert's answer
2019-05-31T00:40:00-0400

A compound is made up of silver and oxygen, it’s an oxide. Oxygen has a valency 2, so the empiric formula is Ag2Ox.

m(O)=m(Ag2Ox)-m(Ag)

m(O)=1.7734-1.6510=0.1224 g

mol=m/M

mol(Ag):mol(O)=1.7734/108 : 0.1224/16=0.0153 : 0.0077=2 : 1

So, the empiric formula of the compound is Ag2O

Let's check:

W(Ag)=m(Ag)/m(Ag2Ox)

W(Ag)=1.6510/1.7734=0.931

W(Ag)=2*M(Ag)/M(Ag2Ox)

W(Ag)=2*108/2*108+16=216/232=0.931

So, the empiric formula is correct.


mol(Ag2O)=m(Ag2O)/M(Ag2O)

M(Ag2O) = 2*M(Ag)+M(O)

M(Ag2O) = 2*108+16=232 g/mol

mol(Ag2O) = 1.7734/232=0.0076 mol


Answer:

M(Ag2O)=232 g/mol

mol(Ag2O)=0.0076 mol

Ag : O = 2 : 1

empirical formula of the compound Ag2O

M(Ag)=108 g/mol

M(O)=16 g/mol